OpenStudy (anonymous):

A kilobyte is 210. A megabyte is 220 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (anonymous):

@henpen

OpenStudy (anonymous):

That doesn't make sense as a question.

OpenStudy (anonymous):

Theories are not supported by evidence. Theories are just guesses. Theories are based on opinions. Theories are supported by experimental evidence. THESE ARE THE ANSWER CHOICES

OpenStudy (anonymous):

What do you think?

OpenStudy (anonymous):

the last one ?

OpenStudy (anonymous):

Physicsts' theory is that gravity exists. Why do they believe that?

OpenStudy (anonymous):

Yes- that' basically how science works- http://www.youtube.com/watch?v=b240PGCMwV0

OpenStudy (anonymous):

okay so it is the last one

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

The link! Watch the link!

ganeshie8 (ganeshie8):

1 kilibyte = 2^10 byte 1 megabyte = 2^20 byte

OpenStudy (dumbcow):

yes that makes more sense ^^

ganeshie8 (ganeshie8):

to find how many kilobytes in a megabyte, just divide 2^20/2^10 = ?